A 3.00 kg pumpkin hit the ground with 49.0 Joules of kinetic energy. How fast was the pumpkin falling? How far did the pumpkin fall?
1
Expert's answer
2020-11-12T09:17:18-0500
"K=0.5mv^2\\49=0.5(3)v^2\\v=5.72\frac{m}{s}"

"t=\frac{v}{g}=\frac{5.72}{9.8}=0.58\ s"